from a standard deck of 52 cards what are the odds of getting a black 7? _:1, what are the odds against it 1:_ Also what are the odds against getting a 4 and in favor of drawing a 4?
Answer by edjones(7569)

There are 2 black 7s.
Probability is 2/52=1/26
odds against: 25 to 1 ratio of unfavorable to favorable possibilities.
odds in favor: 1 to 25
.
There are 4 fours.
probability is 4/52=1/13
odds against 12-1
odds in favor 1-12
.
Ed